# Momentum dependent quantum Ruelle-Pollicott resonances in translationally invariant many-body systems

### v1

@@ Abstract
We study Ruelle-Pollicott resonances in translationally invariant quantum many-body systems via spectra of momentum-resolved operator propagator on infinite systems. Momentum dependence gives insight into decay of correlation functions, showing that, depending on their symmetries, different correlation functions in general decay with different rates. Focusing on the kicked Ising model the spectrum seems to be typically composed of an annular random matrix like ring whose size we theoretically predict, and few isolated resonances. We identify several interesting regimes, including a mixing regime with a power-law decay of correlation functions. In that regime we also observe a huge difference in time-scales of different correlation functions. An exact expression for the singular values of the operator propagator is conjectured, showing that it becomes singular at a special point.
@@ I Introduction
As is often the case tough, theoretical elegance comes with a price. It is not clear how to in practice do the required “analytical continuation”. Mathematics is heavy, and even in exactly solvable classical cases like the cat map antoniou97 , the Bernoulli map or the baker map Hasegawa92 ; Tasaki93 , one often rather starts by writing the spectral decomposition of $U$ on an appropriate system-dependent functional space. Thereby, the crux of the problem shifts to the question of what is the appropriate functional space.
@@ I.1 Previous work
Exact results on RP resonances are possible for classical hyperbolic maps, like the baker and the Bernoulli map Hasegawa92 ; Tasaki93 , and the cat map antoniou97 . The standard map has been studied in Refs. fishman00 ; venegeroles08 , while the kicked top, including a regime where classical dynamics is mixed, in Ref. weber . A detailed analysis of the multibaker map is in Ref. gaspard92 ; gaspard , while coupled baker maps have been studied in Refs. arul94 ; Haake03 . Rather than working with a propagator one can approach everything from the point of view of spectral analysis of correlation functions, extending frequencies to the complex domain gaspard . Thereby one extracts RP resonances directly from correlations, see Ref. martin for an example. More rigorous treatment and a comparison between classical and quantum cases with noise for maps on a torus can be found in Ref. Nonnenmacher . Another technical trick of how to obtain RP resonances is using trace formulas and zeta functions chaosbook . It has been observed that sometimes there might be issues with convergence of RP resonances for particular choices of coarse-graining blank98 ; shepelyansky ; shudo21 (could perhaps be related to strong non-normality of matrices).
@@ I.2 New results
We approach RP resonances by studying the spectrum of a propagator of operators on an infinite system. Resolving the momentum dependence and truncating it down to operators with support on $r$ consecutive sites, we can discuss decay of any correlation function. To illustrate our results we are going to study the same kicked Ising model and use essentially the same truncation method used in Ref. Prosen , but extend it from the zero momentum subspace studied there. Doing that we will find many new and interesting results. As we work directly in the thermodynamic limit, i.e., study evolution of operators in an infinite system, there is only one limit we need to take, namely, the operator support $r$ to infinity, instead of typically two for other types of coarse-graining or noise approaches to RP resonances. We show that to properly assess dynamical properties of the system it pays off to take into account the momentum dependence of RP resonances because different observables in general decay with different rates. The spectrum of eigenvalues in the KI model consists of a random-matrix like bulk of an annular shape whose radii we analytically predict, and few isolated resonances. Those indicate rich physics: besides identifying expected and known chaotic regime, we also find a mixing non-chaotic regime with a power-law decay of correlations due to a branch cut. This regime is rather puzzling because, despite being away from any integrable point, different observables decay on vastly different timescales. For instance, for a particular choice of parameters magnetization correlations start to decay on a $\sim 10^{4}$ times longer time scale than for instance staggered magnetization or magnetization current correlations. The origin of such disparate timescales is not clear. Using analytical expression for singular values of the truncated propagator we also conjecture a lower bound on RP resonances that depends only on the parameter $\tau$. Singular values also highlight speciality of KI parameters when $\tau$ is an odd multiple of $\pi/4$, for which some singular values become 00 (this point includes an even more special dual unitary KI model DU if one in addition sets $h_{\rm x}=1$).
@@ I.3 Kicked Ising model
where the prefactor $1/L$ ensures the asymptotic independence on the system size $L$ in a finite system, and $k$ is the momentum. The expectation value is an infinite temperature average, $\langle\bullet\rangle:=\frac{1}{2^{L}}{\rm tr}(\bullet)$, however in practice, because our system sizes are large enough, it typically suffices to evaluate it in a single random initial state $\psi$, $\langle\bullet\rangle=\langle\psi|\bullet|\psi\rangle$. Namely, fluctuations of $C(t)$ in a single random state are of the order $\sim 1/2^{L/2}$ (e.g., $\approx 10^{-5}$ at $L=32$). We will also briefly consider local observables.
@@ II Momentum-dependent spectrum of Ruelle-Pollicott resonances
Many-body lattice systems with translational invariance have a conserved momentum $k$. Compared to single-particle systems that do not have any spatial degree of freedom one should therefore expect a one-parameter family of RP resonances labeled by that momentum (and any other additional symmetries that one might have). In this section we are going to describe how to write a truncated operator propagator $M(k)$ in the subspace of momentum $k$. The procedure is a generalization of the truncation in the zero momentum sector used in Ref. Prosen .
@@ IV Mixing regime
Looking at numerically computed autocorrelation functions, Fig. 8, we see that indeed the ones from the $k=0^{+}$ do not seem to decay. This looks similar to what was called (for other parameter values) an intermediate regime in Ref. Prosen07 . On the other hand, inspecting other sectors, for instance the current $J$ (12), which is from $k=0^{-}$, or the staggered magnetization $sZ$ (13) which is from $k=\pi$, we can see a clear decay. Decay in fact looks like a power-law $\sim 1/t^{\alpha}$ with the power $\alpha\approx 3$ over more than $2$ decades in $C(t)$.
@@
We can see that at each finite $r$ the asymptotic decay of $\mathbf{y}^{T}M\mathbf{y}$ is actually exponential, in-line with $|\lambda_{1}|<1$, but the rate of this exponential decay decreases with increasing $r$ (the inset). It therefore looks that $|\lambda_{1}(k=0^{-})|$ will in fact with $r$ slowly converge towards $1$. In Fig. 10 we show an example of a full spectrum of $M(k=0^{-})$ that suggests what is going on: in addition to an RMT-like annular cloud of eigenvalues, whose boundaries are well described by Eq.(18), we can also see a large number of eigenvalues on the real axis. That is similar to what has been recently seen arul24 in channels describing reduced dynamics of coupled standard maps in a regime of no KAM tori but with a stable fixed point. We conjecture that as one increases $r$ the number of real eigenvalues increases, pushing the largest one $|\lambda_{1}|$ towards $1$, while the resulting continuum of eigenvalues, i.e., a branch cut, is responsible for the power-law decay. Note that while this scenario is not yet very clear in the shown spectrum for $r=7$, the power-law decay is quite clear in a directly calculated $C(t)$ for $L=33$.
@@
Considering the above explanation we must also rethink the apparent plateau seen in the $k=0^{+}$ sector (Fig. 8). Namely, it seems more natural that one should see decay in all sectors. And indeed, looking at the autocorrelation function of $Z$ (20) at much longer times we eventually do see a decay, Fig. 11. This decay seems to be asymptotically exponential but with an eigenvalue that is very close to $1$, namely $\lambda_{1}\approx 1-15\cdot 10^{-6}$. This exponential decay also starts after very long transient plateau of duration $\sim 10^{3}$. We therefore have a rather interesting situation: observables from the $k=0^{-}$ sector decay as a power law, while those from the $k=0^{+}$ decay exponentially. If this two decays really stay as they are also in the thermodynamic limit (we see no numerical indication that they would not), then mathematically speaking the decay in the $k=0^{+}$ sector would eventually become faster (exponential decay always wins over a power law) than the one in $k=0^{-}$, meaning that the resonance in $k=0^{+}$ is not always the larger. Note that due to prefactors ($0.25\cdot 0.999985^{t}$ vs. $150/t^{3}$) $C(t)$ for the current $J$ would be larger than the one for $Z$ only for $t>2.5\cdot 10^{6}$ when $C(t)<10^{-17}$ (requiring $L>120$ to see it in a single random state). Another possibility is that for larger $L$, when the branch cut in the sector $k=0^{-}$ moves really close to $1$, the largest eigenvalue $\lambda_{1}$ in the sector $k=0^{+}$ is also “pushed” to $1$, at the end also resulting in a power-law decay in the sector $k=0^{+}$.
Inset in Fig. 11, where we show dependence on $\tau$ at a fixed $L$ is also very interesting. It shows that there is a 2nd finite-size plateau at times that rapidly increase as on decreases $\tau$. When one increases $L$ that plateau decreases, as is nicely visible for $\tau=0.45$ in the main plot. Understanding where this large separation of timescales between the even and odd sectors in $k=0$ comes from is a very interesting open problem. Note that there is no obvious small (or large) parameter that could result in this large timescale. If we speculate a bit, it could be that in a many body quantum system upon breaking of integrability one generically immediately becomes ergodic and mixing in the thermodynamic limit, but the mixing timescale could be very large (perhaps akin to Arnold diffusion in classical systems). This must be contrasted with consequences of the KAM theorem in classical systems, for instance, that upon weak perturbations some classical orbits remain essentially integrable, e.g., display non-decaying autocorrelation functions. The fuzziness of quantum mechanics could on a long timescale wash out any possible “KAM tori” that one might have in classical systems.
@@
Figure 11: Autocorrelation function of magnetization $Z$ (20) at $h_{\rm x}=0.9,h_{\rm z}=0.8$. Main plot shows convergence with $L$ at $\tau=0.45$ (dashed black curve is $\sim\lambda_{1}^{t}$ with $\lambda_{1}=1-15\cdot 10^{-6})$, while the inset shows dependence on $\tau$ at fixed system size $L=18$.
@@ V Conclusion
While we essentially always find that the largest resonance is from the even $k=0$ sector, it is not clear if there is a proof that this always has to be the case. In fact, a counterexample could be a very interesting parameter regime where it seems that some correlations decay as a power-law due to a branch cut, while other decay exponentially but with a very small decay rate. It is puzzling and unclear where this separation of timescales comes from, and why breaking of integrability at those parameters is so slow for some observables. While it is believed that the kicked Ising model is a generic chaotic toy model, one can wonder if it is nevertheless special in some way. To that end it would be interesting to study Ruelle-Pollicott resonances in other systems, for instance without a commuting interacting part, including in generic quantum circuits.

### v2

@@ Abstract
We study Ruelle-Pollicott resonances in translationally invariant quantum many-body lattice systems via spectra of momentum-resolved operator propagator on infinite systems. Momentum dependence gives insight into decay of correlation functions, showing that, depending on their symmetries, different correlation functions in general decay with different rates. Focusing on the kicked Ising model the spectrum seems to be typically composed of an annular random matrix like ring whose size we theoretically predict, and few isolated resonances. We identify several interesting regimes, including a mixing regime with a power-law decay of correlation functions. In that regime we also observe a huge difference in time-scales of different correlation functions due to an almost conserved operator. An exact expression for the singular values of the operator propagator is conjectured, showing that it becomes singular at a special point.
@@ I Introduction
As is often the case tough, theoretical elegance comes with a price. It is not clear how to in practice do the required “analytical continuation”. Mathematics is heavy, and even in exactly solvable classical cases like the cat map antoniou97 , the Bernoulli map or the baker map Hasegawa92 ; Tasaki93 (see Ref. gaspard95 for differential equations), one often rather starts by writing the spectral decomposition of $U$ on an appropriate system-dependent functional space. Thereby, the crux of the problem shifts to the question of what is the appropriate functional space.
@@ I.1 Previous work
Exact results on RP resonances are possible for classical hyperbolic maps, like the baker and the Bernoulli map Hasegawa92 ; Tasaki93 , and the cat map antoniou97 . The standard map has been studied in Refs. fishman00 ; venegeroles08 ; mata23 , while the kicked top, including a regime where classical dynamics is mixed, in Ref. weber . A detailed analysis of the multibaker map is in Ref. gaspard92 ; gaspard , while coupled baker maps have been studied in Refs. arul94 ; Haake03 . Rather than working with a propagator one can approach everything from the point of view of spectral analysis of correlation functions, extending frequencies to the complex domain gaspard . Thereby one extracts RP resonances directly from correlations, see Ref. martin for an example. More rigorous treatment and a comparison between classical and quantum cases with noise for maps on a torus can be found in Ref. Nonnenmacher . Another technical trick of how to obtain RP resonances is using periodic orbits with trace formulas and zeta functions chaosbook ; sridhar ; altshuler ; leboeuf . It has been observed that sometimes there might be issues with convergence of RP resonances for particular choices of coarse-graining blank98 ; shepelyansky ; shudo21 (could perhaps be related to strong non-normality of matrices).
@@ I.2 New results
We approach RP resonances by studying the spectrum of a propagator of operators on an infinite system. Resolving the (quasi)momentum dependence and truncating it down to operators with support on $r$ consecutive sites, we can discuss decay of any correlation function. To illustrate our results we are going to study the same kicked Ising model and use essentially the same truncation method used in Ref. Prosen , but extend it from the zero momentum subspace studied there. Doing that we will find many new and interesting results. As we work directly in the thermodynamic limit, i.e., study evolution of operators in an infinite system, there is only one limit we need to take, namely, the operator support $r$ to infinity, instead of typically two for other types of coarse-graining or noise approaches to RP resonances. We show that to properly assess dynamical properties of the system it pays off to take into account the momentum dependence of RP resonances because different observables in general decay with different rates. The spectrum of eigenvalues in the KI model consists of a random-matrix like bulk of an annular shape whose radii we analytically predict, and few isolated resonances. Those indicate rich physics: besides identifying expected and known chaotic regime, we also find a mixing non-chaotic regime with a power-law decay of correlations due to a branch cut. This regime is rather interesting because, despite being away from any integrable point, different observables decay on vastly different timescales. For instance, for a particular choice of parameters magnetization correlations start to decay on a $\sim 10^{4}$ times longer time scale than for instance the staggered magnetization or magnetization current correlations. The origin of such disparate timescales is due to an almost conserved operators giving rise to a prethermalization plateau. Using analytical expression for singular values of the truncated propagator we also conjecture a lower bound on RP resonances that depends only on the parameter $\tau$. Singular values also highlight speciality of KI parameters when $\tau$ is an odd multiple of $\pi/4$, for which some singular values become 00 (this point includes an even more special dual unitary KI model DU if one in addition sets $h_{\rm x}=1$).
@@ I.3 Kicked Ising model
where the prefactor $1/L$ ensures the asymptotic independence on the system size $L$ in a finite system, and $k$ is the (quasi)momentum. The expectation value is an infinite temperature average, $\langle\bullet\rangle:=\frac{1}{2^{L}}{\rm tr}(\bullet)$, however in practice, because our system sizes are large enough, it typically suffices to evaluate it in a single random initial state $\psi$, $\langle\bullet\rangle=\langle\psi|\bullet|\psi\rangle$. Namely, fluctuations of $C(t)$ in a single random state are of the order $\sim 1/2^{L/2}$ (e.g., $\approx 10^{-5}$ at $L=32$). We will also briefly consider local observables.
@@ II Momentum-dependent spectrum of Ruelle-Pollicott resonances
Many-body lattice systems with translational invariance have a conserved quasi-momentum $k$ (we will simply call it momentum). Compared to single-particle systems that do not have any spatial degree of freedom one should therefore expect a one-parameter family of RP resonances labeled by that momentum (and any other additional symmetries that one might have). In this section we are going to describe how to write a truncated operator propagator $M(k)$ in the subspace of momentum $k$. The procedure is a generalization of the truncation in the zero momentum sector used in Ref. Prosen .
@@ IV Mixing regime
IV.1 Power-law decay
Looking at numerically computed autocorrelation functions, Fig. 8, we see that indeed the ones from the $k=0^{+}$ do not seem to decay. This looks similar to what was called (for other parameter values) an intermediate regime in Ref. Prosen07 . On the other hand, inspecting other sectors, for instance the current $J$ (12), which is from $k=0^{-}$, or the staggered magnetization $sZ$ (13) which is from $k=\pi$, we can see a clear decay. Decay in fact looks like a power-law $\sim 1/t^{\alpha}$ with a power $\alpha\approx 3$ over more than $2$ decades in $C(t)$.
@@ IV.1 Power-law decay
We can see that at each finite $r$ the asymptotic decay of $\mathbf{y}^{T}M\mathbf{y}$ is actually exponential, in-line with $|\lambda_{1}|<1$, but the rate of this exponential decay decreases with increasing $r$ (the inset). Increasing $r$ the approximation with $M(0^{-})$ gets better, but the convergence towards numerically computed power-law decay of $C(t)$ for $L=33$ is rather slow. With increasing $r$ the largest eigenvalue $|\lambda_{1}(k=0^{-})|$ will converge towards $1$. In Fig. 10 we show an example of a full spectrum of $M(k=0^{-})$ that suggests what is going on: in addition to an RMT-like annular cloud of eigenvalues, whose boundaries are well described by Eq.(18), we can also see a large number of eigenvalues on the real axis. That is similar to what has been recently seen arul24 in channels describing reduced dynamics of coupled standard maps in a regime of no KAM tori but with a stable fixed point. We conjecture that as one increases $r$ the number of real eigenvalues increases, pushing the largest one $|\lambda_{1}|$ towards $1$, while the resulting continuum of eigenvalues, i.e., a branch cut, is responsible for the power-law decay. Note that while this scenario is not yet very clear in the shown spectrum for $r=7$, the power-law decay is quite clear in a directly calculated $C(t)$ for $L=33$.
@@
IV.2 Prethermalization in $k=0^{+}$
@@
Figure 11: Autocorrelation function of magnetization $Z$ (20) at $h_{\rm x}=0.9,h_{\rm z}=0.8$. Main plot shows convergence with $L$ at $\tau=0.45$ (dashed black curve is $0.25\cdot\lambda_{1}^{t}$ with $\lambda_{1}=1-14\cdot 10^{-6})$, while the inset shows dependence on $\tau$ at fixed system size $L=18$.
Considering the above explanation we must also rethink the apparent plateau seen in the $k=0^{+}$ sector (Fig. 8). Namely, it seems more natural that one would have a decay in all sectors. And indeed, looking at the autocorrelation function of $Z$ (20) at much longer times we eventually do see a decay, Fig. 11. At a finite $L$ there appears to be a 2nd plateau at a late time, e.g. $t>10^{5}$ at $L=20$, which though, as seen in the figure, is just a finite-size effect. As one increases $L$ one approaches the asymptotic exponential decay given by the largest eigenvalue that is very close to $1$, namely $\lambda_{1}\approx 1-14\cdot 10^{-6}$ (for studied truncations $\lambda_{1}$ seems rather stable with only small even-odd $r$ oscillations). This exponential decay only starts after very long transient plateau of duration $\sim 10^{3}$. Looking at the $\tau$ dependence in the inset of Fig. 11 we see that the time and the height of this 2nd finite-size plateau rapidly change with $\tau$. We therefore have a rather interesting situation: observables from the $k=0^{-}$ sector decay (quickly) as a power law, while those from the $k=0^{+}$ eventually do decay exponentially but only after a long plateau.
The long plateau can be explained by what has been called a prethermalization keiji16 . Namely, although $\tau$ does not necessarily look perturbatively small one can nevertheless try to write the effective one-step Floquet operator $H_{\rm eff}$ using a Baker-Campbell-Hausdorff (BCH) expansion
$U={\rm e}^{H_{\rm eff}},\quad H_{\rm eff}$ $=$ $-{\rm i}\tau(H_{\rm x}+H_{\rm z})-\frac{\tau^{2}}{2}[H_{\rm x},H_{\rm z}]+\cdots$ (25)
$=$ $\sum_{p=1}H_{p},$
with $H_{p}$ being the antihermitian $p$-th order term (in $\tau$) in the BCH expansion. If this expansion would converge, and therefore could describe the long-time dynamics by simple $U^{t}={\rm e}^{H_{\rm eff}t}$, the $H_{\rm eff}$ would be a constant of motion. Provided it would be a local operator this should be reflected in the spectral properties of $M(k=0^{+})$, i.e., one eigenvalue being $1$. However, the formal BCH expansion is usually divergent with terms $H_{p}$ eventually growing for $p>p_{0}$. Up until times when divergent $H_{p>p_{0}}$ become important the $H_{\rm eff}$ truncated at the order $p_{0}$ is almost conserved keiji16 , i.e., its exact lack of conservaton will become noticeable only at large times. Such an almost conserved $H_{\rm eff}$, a hallmark of prethermalization keiji16 , will result in one eigenvalue of $M(k=0^{+})$ being very close to $1$. And this is exactly what happens in our kicked Ising model at $\tau=0.45$. This can be seen in Fig. 12. As already mentioned, the largest eigenvalue $\lambda_{1}$ is very close to $1$ with the corresponding eigenvector $\mathbf{x}_{1}$ being, upto normalization, exactly equal to the truncated $H_{\rm eff}$. This is verified in the bottom frame of Fig. 12 where we calculate the overlaps between $\mathbf{x}_{1}$ calculated from $M(0^{+})$ for a finite support size $r$, say $r=3$, and compare it with $H_{\rm eff}$ obtained by summing up the first $p$ orders of the BCH series. Due to the structure of the kicked Ising model the support of the BCH term $H_{p}$ of order $p$ is $\sim 1+p/2$, i.e., every 2nd order increases the support of commutators by $1$ site. This can be nicely seen from the data: for $r=3$ the fidelity saturates at $p\approx 6$, while at $r=5$ at around $p\approx 10$. One can also see that the error of $\mathbf{x}_{1}$ for a given $r$, i.e., its difference to $H_{\rm eff}$, is essentially equal to the norm $|H_{p}|^{2}:={\rm tr}(H_{p}^{\dagger}H_{p})$ of the truncated terms (red diamonds in the figure) for $p\approx 2r$. This all works as long as the BCH series converges, i.e., $H_{p}$ decrease. Namely, by calculating BCH terms upto order $20$ casas08 we can see (Fig. 12) that the terms indeed start to increase after $p_{0}\approx 16$. This means that $H_{\rm eff}$ is not really conserved and we do not expect any eigenvalue of $M(k=0^{+})$ to be exactly $1$. The eventual divergent nature of the BCH series is also reflected in the structure of the eigenvector $\mathbf{x}_{1}$ – for larger $r$ it ceases to be well approximated by $H_{\rm eff}$ (compare $r=5$ and $r=7$ in the figure) – as well as in the spectrum of $M(0^{+})$ (other eigenvalues moving closer to $\lambda_{1}$).
One question that remains is what happens in the thermodynamic limit $L\to\infty$, in particular in view of a seeming power-law decay in other sectors? There are two possibilities. As we can see in the top Fig. 12 there are many eigenvalues on the real axis, the number of which increases with $r$, similarly as is the case in the $k=0^{-}$ sector. So it could be that at very large $r$ when those eigenvalues move really close to $1$ they begin to mix with $\lambda_{1}$ which eventually looses its “prethermalization” character of describing $H_{\rm eff}$, altogether resulting in a power-law decay of correlations also in the $k=0^{+}$ sector. Another, perhaps less likely scenario would instead be that the decays in both $k=0^{+}$ and $k=0^{-}$ stay as they are, an exponential and a power-law, respectively. In such a case the decay in the $k=0^{+}$ sector would eventually become faster than the one in $k=0^{-}$ (exponential decay always wins over a power law), meaning that the resonance in $k=0^{+}$ is not always the largest. Note that due to prefactors ($0.25\cdot 0.999986^{t}$ vs. $150/t^{3}$) $C(t)$ for the current $J$ would be larger than the one for $Z$ only for $t>2.5\cdot 10^{6}$ when $C(t)<10^{-17}$ (requiring $L>120$ to see it in a single random state).
We find it quite interesting that one can extract all prethermalization information from the truncated operator propagator $M(k)$ – a procedure that seems simpler than for instance tediously evaluating BCH terms to a high order. It would be also interesting to understand physical mechanism behind this slow relaxation and whether behavior in different momentum sectors are somehow related. The slow breaking of integrability that we observes is a bit akin to what one expects in classical systems based on the KAM theorem. Upon weak perturbation some classical orbits (KAM torii) remain essentially integrable, e.g., display non-decaying autocorrelation functions, however in a high dimensional system they do not separate the phase space into inside and outside and orbits can escape (Arnold diffusion).
Refer to captionRefer to caption
Figure 12: An almost conserved operator in the mixing regime in sector $k=0^{+}$ for $\tau=0.45,\,h_{\rm x}=0.9,h_{\rm z}=0.8$. Top: Zoom of the spectrum around the largest $\lambda_{1}\approx 1-14\cdot 10^{-6}$ highlighted in red (blue circles are theoretical $r_{\rm in}$ and $r_{\rm out}$). Left is for $r=5$ (423 eigenvalues), right for $r=7$ (6303 eigenvalues). Bottom: The eigenvector $\mathbf{x}_{1}$ corresponding to $\lambda_{1}$ is for small $r$ equal to $H_{\rm eff}$ obtained by summing first $p$ BCH orders. Norm of the BCH term $H_{p}$ of order $p$ is shown with red diamonds.
@@ V Conclusion
While we essentially always find that the largest resonance is from the even $k=0$ sector, it is not clear if there is a proof that this always has to be the case. In fact, a counterexample could be a very interesting parameter regime where it seems that some correlations decay as a power-law due to a branch cut, while other decay exponentially but with a very small decay rate. This separation of timescales can be explained in terms of an almost conserved operator (prethermalization) that can be alternatively calculated via a BCH formula. While it is believed that the kicked Ising model is a generic chaotic toy model we see that its dynamics has a rich dependence on parameters. To that end it would be interesting to study Ruelle-Pollicott resonances in other systems, for instance without a commuting interacting part, including in generic quantum circuits.